# Strong time-periodic solutions to a multilayered fluid-structure interaction problem

### v1

@@ 4.5. Spectral theory
In order to prepare for the following considerations, we collect and discuss several preliminary results in the sequel. Let us start with spectral theory for the fluid-structure operator $A_{\mathrm{fs}}$ as recalled in (4.29) and (4.30). The following result on the spectrum of $A_{\mathrm{fs}}$ can be found in the proof of [41, Thm. 4.6].
[…]
@@
(4.35) $\lambda u-\operatorname{div}\sigma_{\mathrm{f}}(u,\pi)&=0,&&\enspace\text{in}\enspace\Omega_{\mathrm{f}},\\ \operatorname{div}u&=0,&&\enspace\text{in}\enspace\Omega_{\mathrm{f}},\\ \lambda\eta_{1}-\eta_{2}&=0,&&\enspace\text{in}\enspace\omega,\\ \lambda\eta_{2}+P_{\mathrm{m}}\Delta_{s}^{2}\eta_{1}-\Delta_{s}\eta_{2}&=\gamma_{\mathrm{m}}\pi+P_{\mathrm{m}}\bigl{(}\left.(\sigma_{\mathrm{s}}(d_{1},d_{2})\mathrm{e}_{3})\right|_{\Gamma_{\mathrm{s}}}\cdot\mathrm{e}_{3}\bigr{)},&&\enspace\text{in}\enspace\omega,\\ \int_{\omega}\eta_{1}\,\mathrm{d}s&=0,\\ \lambda d_{1}-d_{2}&=0,&&\enspace\text{in}\enspace\Omega_{\mathrm{s}},\\ \lambda d_{2}-\operatorname{div}\sigma_{\mathrm{s}}(d_{1},d_{2})&=0,&&\enspace\text{in}\enspace\Omega_{\mathrm{s}},\\ u&=P_{\mathrm{m}}(\eta_{2})\mathrm{e}_{3},&&\enspace\text{on}\enspace\partial\Omega_{\mathrm{f}}\setminus\Gamma_{0},\\ d_{1}&=P_{\mathrm{m}}(\eta_{1})\mathrm{e}_{3},&&\enspace\text{on}\enspace\Gamma_{\mathrm{s}},\\ u&=0,&&\enspace\text{on}\enspace\Gamma_{0},\\ d_{1}&=0,&&\enspace\text{on}\enspace\partial\Omega_{\mathrm{s}}\setminus\Gamma_{\mathrm{s}},\\ \eta_{1}=\nabla_{s}\eta_{1}\cdot n_{\omega}&=0,&&\enspace\text{on}\enspace\partial\omega.\end{aligned}\right.$
@@
(4.36) $0=\lambda\int_{\Omega_{\mathrm{f}}}|u|^{2}\,\mathrm{d}x+2\int_{\Omega_{\mathrm{f}}}|\mathbb{D}(u)|^{2}\,\mathrm{d}x+\int_{\partial\Omega_{\mathrm{f}}\setminus\Gamma_{0}}\pi\cdot\overline{P_{\mathrm{m}}(\eta_{2})}\,\mathrm{d}s.$
@@
(4.37) $\lambda\int_{\omega}|\eta_{2}|^{2}\,\mathrm{d}s+\overline{\lambda}\int_{\omega}|\Delta_{s}\eta_{1}|^{2}\,\mathrm{d}s+\int_{\omega}|\nabla_{s}\eta_{2}|^{2}\,\mathrm{d}s=\int_{\omega}\gamma_{\mathrm{m}}\pi\overline{\eta}_{2}\,\mathrm{d}s+\int_{\omega}P_{\mathrm{m}}\left.\bigl{(}\sigma_{\mathrm{s}}(d_{1},d_{2})\mathrm{e}\bigr{)}\right|_{\Gamma_{\mathrm{s}}}\cdot\mathrm{e}_{3}\overline{\eta}_{2}\,\mathrm{d}s.$
@@
$\quad+\int_{\omega}|\nabla_{s}\eta_{2}|^{2}\,\mathrm{d}s-\int_{\omega}P_{\mathrm{m}}\left.\bigl{(}\sigma_{\mathrm{s}}(d_{1},d_{2})\mathrm{e}\bigr{)}\right|_{\Gamma_{\mathrm{s}}}\cdot\mathrm{e}_{3}\overline{\eta}_{2}\,\mathrm{d}s=0.$
@@
$\quad-\overline{\lambda}\int_{\partial\Omega_{\mathrm{s}}}S_{\mathrm{s}}(d_{1})\overline{d}_{1}\cdot n\,\mathrm{d}x-\delta\int_{\partial\Omega_{\mathrm{s}}}S_{\mathrm{s}}(d_{2})\overline{d}_{2}\cdot n\,\mathrm{d}x=-\int_{\Gamma_{\mathrm{s}}}S_{\mathrm{s}}(d_{1})\overline{d}_{2}\cdot n\,\mathrm{d}s-\delta\int_{\Gamma_{\mathrm{s}}}S_{\mathrm{s}}(d_{2})\overline{d}_{2}\cdot n\,\mathrm{d}s$
$=-\overline{\lambda}\int_{\Gamma_{\mathrm{s}}}S_{\mathrm{s}}(d_{1}+\delta d_{2})\overline{d}_{1}\cdot n\,\mathrm{d}s=-\overline{\lambda}\int_{\Gamma_{\mathrm{s}}}\sigma_{\mathrm{s}}(d_{1},d_{2})\overline{P_{\mathrm{m}}(\eta_{1})\mathrm{e}_{3}}\,\mathrm{d}s=-\int_{\Gamma_{\mathrm{s}}}\sigma_{\mathrm{s}}(d_{1},d_{2})P_{\mathrm{m}}(\overline{\eta}_{2})\mathrm{e}_{3}\,\mathrm{d}s.$
@@
00 0 $=\lambda\int_{\Omega_{\mathrm{f}}}|u|^{2}\,\mathrm{d}x+2\int_{\Omega_{\mathrm{f}}}|\mathbb{D}(u)|^{2}\,\mathrm{d}x+\lambda\int_{\omega}|\eta_{2}|^{2}\,\mathrm{d}s+\overline{\lambda}\int_{\omega}|\Delta_{s}\eta_{1}|^{2}\,\mathrm{d}s+\int_{\omega}|\nabla_{s}\eta_{2}|^{2}\,\mathrm{d}s+\lambda\int_{\Omega_{\mathrm{s}}}|d_{2}|^{2}\,\mathrm{d}x$
@@
We now successively study the spectrum of the operator $\tilde{A}_{\mathrm{mfs}}$. To this end, let us begin with the diagonal operator matrix $\operatorname{diag}(A_{\mathrm{fs}},A_{\mathrm{s}})$ with domain $\mathrm{D}(\tilde{A}_{\mathrm{mfs}})=\mathrm{D}(A_{\mathrm{fs}})\times\mathrm{D}(A_{\mathrm{s}})$. From Subsection 4.5 together with Subsection 4.2, it first follows that $\sigma(\operatorname{diag}(A_{\mathrm{fs}},A_{\mathrm{s}}))\subset\{\lambda\in\mathbb{C}:\operatorname{Re}\lambda<0\}$, so $s(\operatorname{diag}(A_{\mathrm{fs}},A_{\mathrm{s}}))<0$. Next, we recall from (4.33) and (4.34) the operator matrix $\tilde{A}_{\mathrm{mfs}}^{\mathrm{t}}\colon\mathrm{D}(\tilde{A}_{\mathrm{mfs}}^{\mathrm{t}})\subset\mathrm{Y}_{0}\to\mathrm{Y}_{0}$ and the representation
$\lambda-\tilde{A}_{\mathrm{mfs}}^{\mathrm{t}}=\begin{pmatrix}\lambda-A_{\mathrm{fs}}&0\\ -\tilde{C}&\lambda-A_{\mathrm{s}}\end{pmatrix}=\begin{pmatrix}\operatorname{Id}&0\\ -\tilde{C}R(\lambda,A_{\mathrm{fs}})&\operatorname{Id}\end{pmatrix}\cdot\begin{pmatrix}\lambda-A_{\mathrm{fs}}&0\\ 0&\lambda-A_{\mathrm{s}}\end{pmatrix},\enspace\text{for}\enspace\lambda\in\rho(A_{\mathrm{fs}}).$
For $\lambda\in\rho(\operatorname{diag}(A_{\mathrm{fs}},A_{\mathrm{s}}))$, we find $\lambda\in\rho(\tilde{A}_{\mathrm{mfs}}^{\mathrm{t}})$, so $\sigma(\tilde{A}_{\mathrm{mfs}}^{\mathrm{t}})\subset\{\lambda\in\mathbb{C}:\operatorname{Re}\lambda<0\}$ and $s(\tilde{A}_{\mathrm{mfs}}^{\mathrm{t}})<0$.
With regard to (4.32), it thus remains to discuss $\tilde{B}=\begin{pmatrix}-BT&B\\ 0&TB\end{pmatrix}$. In the proof of Subsection 4.4, we have already verified that $\tilde{B}$ is a relatively compact perturbation of $\tilde{A}_{\mathrm{mfs}}^{\mathrm{t}}$. From [30, Thm. 5.35] and the above considerations, we infer that
$\sigma_{\mathrm{ess}}(\tilde{A}_{\mathrm{mfs}})=\sigma_{\mathrm{ess}}(\tilde{A}_{\mathrm{mfs}}^{\mathrm{t}})\subset\sigma(\tilde{A}_{\mathrm{mfs}}^{\mathrm{t}})\subset\{\lambda\in\mathbb{C}:\operatorname{Re}\lambda<0\}.$
@@
The above proposition also allows us to deduce that $-A_{\mathrm{mfs}}$ admits maximal $\mathrm{L}^{p}$-regularity on $\mathrm{X}_{0}$ without shift, and that the associated semigroup is exponentially stable.
[…]
Let $q\in(1,\infty)$, and recall the multilayered fluid-structure operator $A_{\mathrm{mfs}}$ from (4.25). Then $-A_{\mathrm{mfs}}$ has maximal $\mathrm{L}^{p}$-regularity, and the associated semigroup $\mathrm{e}^{tA_{\mathrm{mfs}}}$ is exponentially stable, i. e., there exist $\mu_{0}>0$ and $C>0$ such that $\|\mathrm{e}^{tA_{\mathrm{mfs}}}w\|_{\mathrm{X}_{0}}\leq C\mathrm{e}^{-\mu_{0}t}\cdot\|w\|_{\mathrm{X}_{0}}$ for all $w\in\mathrm{X}_{0}$.
@@ 4.6. Maximal periodic regularity of the multilayered fluid-structure operator
Let $q\in(1,\infty)$, and consider the operator $A_{\mathrm{mfs}}$ as introduced in (4.25). Then $-A_{\mathrm{mfs}}$ has maximal periodic $\mathrm{L}^{p}$-regularity on $\mathrm{X}_{0}$.
@@
By construction, it readily follows that $\mathbb{P}f\in\mathrm{L}^{p}(0,T;\mathrm{L}_{\sigma}^{q}(\Omega_{\mathrm{f}}))$. Therefore, together with the preceding argument, we conclude that $(\mathbb{P}f,0,\overline{g},0,h)\in\mathrm{L}^{p}(0,T;\mathrm{X}_{0})$, where $\mathrm{X}_{0}$ is as defined in (4.24). For the domain $\mathrm{D}(A_{\mathrm{mfs}})$ of the multilayered fluid-structure operator as introduced in (4.26), it then follows that there exists a unique solution $(\mathbb{P}u,\eta_{1},\eta_{2},d_{1},d_{2})\in\mathrm{W}^{1,p}(0,T;\mathrm{X}_{0})\cap\mathrm{L}^{p}(0,T;\mathrm{D}(A_{\mathrm{mfs}}))$ to the time-periodic analogue of (4.27). It remains to argue that $u$ and $\pi$ possess the desired regularities. To this end, from Subsection 4.4, we recall $(\operatorname{Id}-\mathbb{P})u=(\operatorname{Id}-\mathbb{P})D_{\mathrm{fl}}\eta_{2}$, where $D_{\mathrm{fl}}$ has been introduced in (4.12). Moreover, from Subsection 4.3, we recall that $D_{\mathrm{fl}}\in\mathcal{L}(\mathrm{W}_{0}^{2,q}(\omega)\cap\mathrm{L}_{\mathrm{m}}^{q}(\omega),\mathrm{W}^{2,q}(\Omega_{\mathrm{f}})^{3})$. On the other hand, arguing as in [41, (4.27)], we get $D_{\mathrm{fl}}\in\mathcal{L}(\mathrm{L}_{\mathrm{m}}^{q}(\omega),\mathrm{L}^{q}(\Omega_{\mathrm{f}})^{3})$. In total, this reveals that $u\in\mathbb{E}_{1}^{u}$.
For the pressure, we invoke $\pi=N(\Delta\mathbb{P}u\cdot\nu)-N_{1}\eta_{2}^{\prime}+N_{2}f$ from Subsection 4.4. From (4.15)1 and (4.15)2, we get $N(\Delta\mathbb{P}u\cdot\nu)\in\mathrm{L}^{p}(0,T;\mathrm{W}^{1,q}(\Omega_{\mathrm{f}})\cap\mathrm{L}_{\mathrm{m}}^{q}(\Omega_{\mathrm{f}}))$ and $N_{1}\eta_{2}^{\prime}\in\mathrm{L}^{p}(0,T;\mathrm{W}^{1,q}(\Omega_{\mathrm{f}})\cap\mathrm{L}_{\mathrm{m}}^{q}(\Omega_{\mathrm{f}}))$, respectively, while the mapping properties of $N_{2}$ from (4.18) yield that we also have $N_{2}f\in\mathrm{L}^{p}(0,T;\mathrm{W}^{1,q}(\Omega_{\mathrm{f}})\cap\mathrm{L}_{\mathrm{m}}^{q}(\Omega_{\mathrm{f}}))$. In summary, we have verified that $\pi\in\mathbb{E}_{1}^{\pi}$. ∎

### v2

@@ 5.5. Spectral theory
In order to prepare for the following considerations, we collect and discuss several preliminary results in the sequel. Let us start with spectral theory for the fluid-structure operator $A_{\mathrm{fs}}$ as recalled in (5.31) and (5.32). The following result on the spectrum of $A_{\mathrm{fs}}$ can be found in the proof of [48, Thm. 4.6].
Lemma 5.15.
@@
(5.37) $\left\{\begin{aligned} \lambda u-\operatorname{div}\sigma_{\mathrm{f}}(u,\pi)&=0,&&\kern 5.0pt\text{in}\kern 5.0pt\Omega_{\mathrm{f}},\\ \operatorname{div}u&=0,&&\kern 5.0pt\text{in}\kern 5.0pt\Omega_{\mathrm{f}},\\ \lambda\eta_{1}-\eta_{2}&=0,&&\kern 5.0pt\text{in}\kern 5.0pt\omega,\\ \lambda\eta_{2}+P_{\mathrm{m}}\Delta_{s}^{2}\eta_{1}-\Delta_{s}\eta_{2}&=\gamma_{\mathrm{m}}\pi+P_{\mathrm{m}}\bigl(\left.(\sigma_{\mathrm{s}}(d_{1},d_{2})\mathrm{e}_{3})\right|_{\Gamma_{0}}\cdot\mathrm{e}_{3}\bigr),&&\kern 5.0pt\text{in}\kern 5.0pt\omega,\\ \int_{\omega}\eta_{1}\,\mathrm{d}s&=0,\\ \lambda d_{1}-d_{2}&=0,&&\kern 5.0pt\text{in}\kern 5.0pt\Omega_{\mathrm{s}},\\ \lambda d_{2}-\operatorname{div}\sigma_{\mathrm{s}}(d_{1},d_{2})&=0,&&\kern 5.0pt\text{in}\kern 5.0pt\Omega_{\mathrm{s}},\\ u&=P_{\mathrm{m}}(\eta_{2})\mathrm{e}_{3},&&\kern 5.0pt\text{on}\kern 5.0pt\Gamma_{0},\\ d_{1}&=P_{\mathrm{m}}(\eta_{1})\mathrm{e}_{3},&&\kern 5.0pt\text{on}\kern 5.0pt\Gamma_{0},\\ u&=0,&&\kern 5.0pt\text{on}\kern 5.0pt\Gamma_{\mathrm{f}},\\ d_{1}&=0,&&\kern 5.0pt\text{on}\kern 5.0pt\Gamma_{\mathrm{s}},\\ \eta_{1}=\nabla_{s}\eta_{1}\cdot n_{\omega}&=0,&&\kern 5.0pt\text{on}\kern 5.0pt\partial\omega.\end{aligned}\right$.
@@
(5.38) $0=\lambda\int_{\Omega_{\mathrm{f}}}|u|^{2}\,\mathrm{d}x+2\int_{\Omega_{\mathrm{f}}}|\mathbb{D}(u)|^{2}\,\mathrm{d}x+\int_{\Gamma_{0}}\pi\cdot\overline{P_{\mathrm{m}}(\eta_{2})}\,\mathrm{d}s.$
@@
(5.39) $\lambda\int_{\omega}|\eta_{2}|^{2}\,\mathrm{d}s+\overline{\lambda}\int_{\omega}|\Delta_{s}\eta_{1}|^{2}\,\mathrm{d}s+\int_{\omega}|\nabla_{s}\eta_{2}|^{2}\,\mathrm{d}s=\int_{\omega}\gamma_{\mathrm{m}}\pi\overline{\eta}_{2}\,\mathrm{d}s+\int_{\omega}P_{\mathrm{m}}\left.\bigl(\sigma_{\mathrm{s}}(d_{1},d_{2})\mathrm{e}\bigr)\right|_{\Gamma_{0}}\cdot\mathrm{e}_{3}\overline{\eta}_{2}\,\mathrm{d}s.$
@@
$\quad+\int_{\omega}|\nabla_{s}\eta_{2}|^{2}\,\mathrm{d}s-\int_{\omega}P_{\mathrm{m}}\left.\bigl(\sigma_{\mathrm{s}}(d_{1},d_{2})\mathrm{e}_{3}\bigr)\right|_{\Gamma_{0}}\cdot\mathrm{e}_{3}\overline{\eta}_{2}\,\mathrm{d}s=0.$
@@
$\quad-\overline{\lambda}\int_{\partial\Omega_{\mathrm{s}}}S_{\mathrm{s}}(d_{1})\overline{d}_{1}\cdot n\,\mathrm{d}x-\delta\int_{\partial\Omega_{\mathrm{s}}}S_{\mathrm{s}}(d_{2})\overline{d}_{2}\cdot n\,\mathrm{d}x=-\int_{\Gamma_{0}}S_{\mathrm{s}}(d_{1})\overline{d}_{2}\cdot n\,\mathrm{d}s-\delta\int_{\Gamma_{0}}S_{\mathrm{s}}(d_{2})\overline{d}_{2}\cdot n\,\mathrm{d}s$
$=-\overline{\lambda}\int_{\Gamma_{0}}S_{\mathrm{s}}(d_{1}+\delta d_{2})\overline{d}_{1}\cdot n\,\mathrm{d}s=-\overline{\lambda}\int_{\Gamma_{0}}\sigma_{\mathrm{s}}(d_{1},d_{2})\overline{P_{\mathrm{m}}(\eta_{1})\mathrm{e}_{3}}\,\mathrm{d}s=-\int_{\Gamma_{0}}\sigma_{\mathrm{s}}(d_{1},d_{2})P_{\mathrm{m}}(\overline{\eta}_{2})\mathrm{e}_{3}\,\mathrm{d}s.$
@@
$0$ $=\lambda\int_{\Omega_{\mathrm{f}}}|u|^{2}\,\mathrm{d}x+2\int_{\Omega_{\mathrm{f}}}|\mathbb{D}(u)|^{2}\,\mathrm{d}x+\lambda\int_{\omega}|\eta_{2}|^{2}\,\mathrm{d}s+\overline{\lambda}\int_{\omega}|\Delta_{s}\eta_{1}|^{2}\,\mathrm{d}s+\int_{\omega}|\nabla_{s}\eta_{2}|^{2}\,\mathrm{d}s+\lambda\int_{\Omega_{\mathrm{s}}}|d_{2}|^{2}\,\mathrm{d}x$
@@
We now successively study the spectrum of the operator $\tilde{A}_{\mathrm{mfs}}$. To this end, let us begin with the diagonal operator matrix $\operatorname{diag}(A_{\mathrm{fs}},A_{\mathrm{s}})$ with domain $\mathrm{D}(\tilde{A}_{\mathrm{mfs}})=\mathrm{D}(A_{\mathrm{fs}})\times\mathrm{D}(A_{\mathrm{s}})$. From Subsection 5.5 together with Subsection 5.2, it first follows that $\sigma(\operatorname{diag}(A_{\mathrm{fs}},A_{\mathrm{s}}))\subset\{\lambda\in\mathbb{C}:\operatorname{Re}\lambda<0\}$, so $s(\operatorname{diag}(A_{\mathrm{fs}},A_{\mathrm{s}}))<0$. Next, we recall from (5.35) and (5.36) the operator matrix $\tilde{A}_{\mathrm{mfs}}^{\triangle}\colon\mathrm{D}(\tilde{A}_{\mathrm{mfs}}^{\triangle})\subset\mathrm{Y}_{0}\to\mathrm{Y}_{0}$ and the representation
$\lambda-\tilde{A}_{\mathrm{mfs}}^{\triangle}=\begin{pmatrix}\lambda-A_{\mathrm{fs}}&0\\ -\tilde{C}&\lambda-A_{\mathrm{s}}\end{pmatrix}=\begin{pmatrix}\operatorname{Id}&0\\ -\tilde{C}R(\lambda,A_{\mathrm{fs}})&\operatorname{Id}\end{pmatrix}\cdot\begin{pmatrix}\lambda-A_{\mathrm{fs}}&0\\ 0&\lambda-A_{\mathrm{s}}\end{pmatrix},\kern 5.0pt\text{for}\kern 5.0pt\lambda\in\rho(A_{\mathrm{fs}}).$
For $\lambda\in\rho(\operatorname{diag}(A_{\mathrm{fs}},A_{\mathrm{s}}))$, we find $\lambda\in\rho(\tilde{A}_{\mathrm{mfs}}^{\triangle})$, so $\sigma(\tilde{A}_{\mathrm{mfs}}^{\triangle})\subset\{\lambda\in\mathbb{C}:\operatorname{Re}\lambda<0\}$ and $s(\tilde{A}_{\mathrm{mfs}}^{\triangle})<0$.
With regard to (5.34), it thus remains to discuss $\tilde{B}=\begin{pmatrix}-BT&B\\ 0&TB\end{pmatrix}$. In the proof of Subsection 5.4, we have already verified that $\tilde{B}$ is a relatively compact perturbation of $\tilde{A}_{\mathrm{mfs}}^{\triangle}$. From [37, Thm. 5.35] and the above considerations, we infer that
$\sigma_{\mathrm{ess}}(\tilde{A}_{\mathrm{mfs}})=\sigma_{\mathrm{ess}}(\tilde{A}_{\mathrm{mfs}}^{\triangle})\subset\sigma(\tilde{A}_{\mathrm{mfs}}^{\triangle})\subset\{\lambda\in\mathbb{C}:\operatorname{Re}\lambda<0\}.$
@@
The above proposition also allows us to deduce that $A_{\mathrm{mfs}}$ admits maximal $\mathrm{L}^{p}$-regularity on $\mathrm{X}_{0}$ without shift, and that the associated semigroup is exponentially stable.
Corollary 5.18.
Let $q\in(1,\infty)$, and recall the multilayered fluid-structure operator $A_{\mathrm{mfs}}$ from (5.27). Then $A_{\mathrm{mfs}}$ has maximal $\mathrm{L}^{p}$-regularity, and the associated semigroup $\mathrm{e}^{tA_{\mathrm{mfs}}}$ is exponentially stable, i.e., there exist $\mu_{0}>0$ and $C>0$ such that $\|\mathrm{e}^{tA_{\mathrm{mfs}}}w\|_{\mathrm{X}_{0}}\leq C\mathrm{e}^{-\mu_{0}t}\cdot\|w\|_{\mathrm{X}_{0}}$ for all $w\in\mathrm{X}_{0}$.
@@ 5.6. Maximal periodic regularity of the multilayered fluid-structure operator
Let $q\in(1,\infty)$, and consider the operator $A_{\mathrm{mfs}}$ as introduced in (5.27). Then $A_{\mathrm{mfs}}$ has maximal periodic $\mathrm{L}^{p}$-regularity on $\mathrm{X}_{0}$.
@@
By construction, it readily follows that $\mathbb{P}f\in\mathrm{L}^{p}(0,T;\mathrm{L}_{\sigma}^{q}(\Omega_{\mathrm{f}}))$. Therefore, together with the preceding argument, we conclude that $(\mathbb{P}f,0,\overline{g},0,h)\in\mathrm{L}^{p}(0,T;\mathrm{X}_{0})$, where $\mathrm{X}_{0}$ is as defined in (5.26). For the domain $\mathrm{D}(A_{\mathrm{mfs}})$ of the multilayered fluid-structure operator as introduced in (5.28), it then follows that there exists a unique solution $(\mathbb{P}u,\eta_{1},\eta_{2},d_{1},d_{2})\in\mathrm{W}^{1,p}(0,T;\mathrm{X}_{0})\cap\mathrm{L}^{p}(0,T;\mathrm{D}(A_{\mathrm{mfs}}))$ to the time-periodic analogue of (5.29). It remains to argue that $u$ and $\pi$ possess the desired regularities. To this end, from Subsection 5.4, we recall $(\operatorname{Id}-\mathbb{P})u=(\operatorname{Id}-\mathbb{P})D_{\mathrm{fl}}\eta_{2}$, where $D_{\mathrm{fl}}$ has been introduced in (5.14). Moreover, from Subsection 5.3, we recall that $D_{\mathrm{fl}}\in\mathcal{L}(\mathrm{W}_{0}^{2,q}(\omega)\cap\mathrm{L}_{\mathrm{m}}^{q}(\omega),\mathrm{W}^{2,q}(\Omega_{\mathrm{f}})^{3})$. On the other hand, arguing as in [48, (4.27)], we get $D_{\mathrm{fl}}\in\mathcal{L}(\mathrm{L}_{\mathrm{m}}^{q}(\omega),\mathrm{L}^{q}(\Omega_{\mathrm{f}})^{3})$. In total, this reveals that $u\in\mathbb{E}_{1}^{u}$.
For the pressure, we invoke $\pi=N(\Delta\mathbb{P}u\cdot\nu)-N_{1}\eta_{2}^{\prime}+N_{2}f$ from Subsection 5.4. From (5.17)1 and (5.17)2, we get $N(\Delta\mathbb{P}u\cdot\nu)\in\mathrm{L}^{p}(0,T;\mathrm{W}^{1,q}(\Omega_{\mathrm{f}})\cap\mathrm{L}_{\mathrm{m}}^{q}(\Omega_{\mathrm{f}}))$ and $N_{1}\eta_{2}^{\prime}\in\mathrm{L}^{p}(0,T;\mathrm{W}^{1,q}(\Omega_{\mathrm{f}})\cap\mathrm{L}_{\mathrm{m}}^{q}(\Omega_{\mathrm{f}}))$, respectively, while the mapping properties of $N_{2}$ from (5.20) yield that $N_{2}f\in\mathrm{L}^{p}(0,T;\mathrm{W}^{1,q}(\Omega_{\mathrm{f}})\cap\mathrm{L}_{\mathrm{m}}^{q}(\Omega_{\mathrm{f}}))$. In summary, we have verified that $\pi\in\mathbb{E}_{1}^{\pi}$. ∎